OpenStudy (samigupta8):

if a,b and c are the angles made by any vector with x,y and z axis respectively then find the value of cos^2a+cos^2b-sin^2c

OpenStudy (anonymous):

As you should know: sin^2c=1-cos^2c Then substituting this inot the equation: cos^2a+cos^2b-(1-cos^2c)=cos^2a+cos^2b+cos^2c-1 Thn another identity: cos^2a+cos^2b+cos^2c=1 So the result is =0

OpenStudy (samigupta8):

but how cos^2a+cos^2b+cos^2c=1

OpenStudy (samigupta8):

hmmm...

OpenStudy (samigupta8):

no i m not gud in drawing

OpenStudy (samigupta8):

can u sort it out 4 me

OpenStudy (loser66):

@terenzreignz help him, please, I don't know

terenzreignz (terenzreignz):

Me neither :3 Me and vectors don't get along T.T @hartnn ?

OpenStudy (samigupta8):

plsssss....anybody just help me out

ganeshie8 (ganeshie8):